Sunshine Media has just completed an initial public offering, where 50 million shares of the 125 million shares outstanding were issued to the public at an offering price of $ 22 per share. On the offering date, the stock price zoomed to $ 40 per share. Who gains from this increase in the price? Who loses, and how much?
Answer #1

Value of shares at offering price = 50,000,000 * 22

Value of shares at offering price = $1,100,000,000

Value of shares at market price = 50,000,000 * 40

Value of shares at market price = $2,000,000,000

Difference = 2,000,000,000 - 1,100,000,000 = $900,000,000

$900,000,000 is the gain to investors who bought shares from Sunshine Media at the offering price

$900,000,000 is the loss to Sunshine Media who sold shares to investors at the offering price